### v1

# Sparse identification of quasipotentials via a combined data-driven method

Abstract
The quasipotential function allows for comprehension and prediction of the escape mechanisms from metastable states in nonlinear dynamical systems. This function acts as a natural extension of the potential function for non-gradient systems and it unveils important properties such as the maximum likelihood transition paths, transition rates and expected exit times of the system. Here, we leverage on machine learning via the combination of two data-driven techniques, namely a neural network and a sparse regression algorithm, to obtain symbolic expressions of quasipotential functions. The key idea is first to determine an orthogonal decomposition of the vector field that governs the underlying dynamics using neural networks, then to interpret symbolically the downhill and circulatory components of the decomposition. These functions are regressed simultaneously with the addition of mathematical constraints. We show that our approach discovers a parsimonious quasipotential equation for an archetypal model with a known exact quasipotential and for the dynamics of a nanomechanical resonator. The analytical forms deliver direct access to the stability of the metastable states and predict rare events with significant computational advantages. Our data-driven approach is of interest for a wide range of applications in which to assess the fluctuating dynamics.
[…]
1 Introduction
Examining the stability of attractors and investigating noise-induced switching between metastable solutions is indispensable for understanding significant biophysical, engineering and ecological phenomena [1, 2], including the folding of proteins [3], the kinetics of chemical reactions [4], and the dynamics of mechanical devices [5, 6]. Comprehensive knowledge of activated escape from a metastable state is crucial for mastering non-equilibrium phenomena that arise from the interplay between nonlinearity and fluctuations [7, 8, 9].
An effective method for illustrating the physics of fluctuation-activated transitions between coexisting stable states is the so-called “ball-in-cup” diagram analogy. In this representation, the state of the dynamical system is depicted by the ball’s position as it rolls across a surface. Random perturbations have the potential to shift the ball from one basin of attraction to another. This surface is defined by a potential function, serving as a natural metric for stability. However, it is worth noting that systems suitable for this depiction, known as gradient systems, are relatively rare [10]. Indeed, when attempting to model the non-equilibrium statistics of transitions between stable states, it is common to encounter non-gradient systems [11]. In this case, the transition dynamics becomes predictable and is entirely defined by the quasipotential that extends the idea of energy functions [12]. As a result, the quasipotential landscape provides an intuitive picture of the fundamental dynamical characteristics of complex systems operating outside of equilibrium [13]. With the quasipotential at hand, valuable asymptotic information can be easily estimated. For instance, a calculation over the energy surface allows for the straightforward determination of the maximum likelihood transition path [14]. Additionally, a precise estimate for the invariant probability distribution near an attractor and the expected escape time from a basin of attraction has been formulated on the basis of the quasipotential topology [15].
Although the quasipotential holds considerable importance for being the informative landscapes in stochastic dynamical systems, calculating it presents a formidable challenge. As a matter of fact, it can be analytically derived only in in some special cases [16]. Ono can rely on various minimum action methods designed to compute the quasipotential between two specified points by solving a path-space minimization problem [17]. Differently, others focus on solving a functional minimization problem, aimed to assess the quasipotential landscapes placing emphasis on the action functional. Nevertheless, all the approaches, such as the variational formulation based on the Freidlin-Wentzell action functional or solving the associated Hamilton-Jacobi equation over a mesh of spatial points, pose difficult numerical challenges with prohibitive computational expense. This is where machine learning (ML) techniques enter overwhelming into the picture. Even if still in early stage of development, data-driven methods and ML-based approaches have significantly impacted the realm of nonlinear dynamics. The emerging scaling of ML techniques have demonstrated capabilities in both interpretation and development of new identification methods [18, 19, 20].
The quasipotential is identified via ML in Lin et al. [21], where an efficient neural network approach is employed to find the orthogonal decomposition of the vector field that governs the dynamics. The idea of utilizing a decomposition of the vector field was also used to compute the generalized potential or invariant distribution for randomly perturbed dynamical systems [22, 23]. In [24], a neural network with automatic differentiation inspired by physics-informed neural network is employed to compute the quasipotential based on the Hamilton-Jacobi equation. A neural network estimation of the quasipotential also promotes for the strategic control of the mean exit time for general stochastic dynamical systems [25].
In this study, we introduce a new data-driven approach to achieve a symbolic interpretation of the quasipotential function. The neural network-based approach of Lin et al. and the sparse regression identification technique developed by [26], are combined to find an analytical model. We leverage the capabilities of neural networks to find the orthogonal decomposition of the vector field driving the dynamics. Subsequently, the functions are reinterpreted using constrained regression within a suitable basis of functions. The quasipotential equation has parsimonious and interpretable form, in the sense that the model is sparse in the space of all the possible functions [27, 28]. In our method, observed trajectories of the dynamical system are the only input required by the combined approach and the goal is to identify a symbolic quasipotential function using a fully data-driven methodology.
The paper is organized as follows. In Sec. 2, we introduce the background of the quasipotential for non-gradient systems and the two parts of the data-driven method. In Sec. 3, two applications of the method are presented. In Sec. 4, we draw the conclusions.
2 Methods
2.1 Expanding on the concept of potential function
We undergo the investigation and interpretation of the process through which a dynamical systems evolve when in presence of multiple equilibra. The underlying physics of transient dynamics and attractors is assumed to be described by the state equation
$\dot{\mathbf{x}}(t)=\mathbf{f}\left(\mathbf{x}(t)\right),$ (1)
where $\dot{\mathbf{x}}:=d\mathbf{x}/dt$ represents the rate of change of the state variables $\mathbf{x}$ over time, and $\mathbf{f}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denotes the nonlinear vector field of dimension $d$ governing the dynamics and responsible for its multistability. In the particular case in which
$\mathbf{f}=-\nabla U(\mathbf{x})$ (2)
we are in the presence of a gradient systems, with $U$ referred to as the “potential”. Thanks to this function, the dynamics of this system can be likened to that of a ball-in-cup setup, where the surface is defined by the potential. In a more general scenario, i.e. non-gradient systems, it is not possible to find $U$ such that Eq. (2) is satisfied. However, the concept of the potential function can be extended to encompass non-gradient systems via the Freidlin-Wentzell quasipotential function [12]. To determine this function, we decompose the vector field $\mathbf{f}$ as
$\mathbf{f}(\mathbf{x})=-\nabla V(\mathbf{x})+\mathbf{g}(\mathbf{x})$ (3)
where $V$ is solution of the Hamilton-Jacobi equation
$\nabla V\cdot\nabla V+\mathbf{f}\cdot\nabla V=0.$ (4)
Under suitable extra conditions for $V$ [12, 21], the function $V$ is locally a scalar multiple of a function known as the Freidlin-Wentzell quasipotential $U$, i.e. $U=2V$. It represents a useful generalization of the potential function for a n-dimensional non-gradient system. In terms of the vector field, the $-\nabla V$ component in Eq. (3) spurs a ball to roll to the bottom of a valley in the ball-in-a-cup analogy, thus $V$ specifies an hyper-surface in which all the trajectories move “downhill” in the absence of perturbations. From Eq. (3) and the Hamilton-Jacobi relation (4) it is straightforward to verify that
$\mathbf{g}\cdot\nabla V=0,$ (5)
hence, $\mathbf{g}$ and $\nabla V$ are perpendicular. Without additional external forces, the “circulatory” component $\mathbf{g}$ creates the circulation of trajectory around levels of $V$. For convenience, we refer to $\nabla V(\mathbf{x})$ and $\mathbf{g}(\mathbf{x})$ in the decomposition (3), (5) as downhill and circulatory components, respectively.
Our method derives $V$ without solving numerically Eq. (4). A two-step data-driven approach is designed to output the symbolic expression of the quasipotential in terms of the state space variables, i.e. $V(\mathbf{x})$. First, a neural network determines the elements in the orthogonal decomposition of Eq. (3) (Sec. 2.2), second, the learned mapping functions are reinterpreted via a constrained regression (Sec. 2.3).
2.2 Dataset and neural networks
The downhill and circulatory components in the orthogonal decomposition (3) are identified with two single neural networks, namely $V_{\theta}$ and $\mathbf{g}_{\theta}$. Hence, the parameterized vector field becomes
$\mathbf{f}_{\theta}(\mathbf{x})=-\nabla V_{\theta}(\mathbf{x})+\mathbf{g}_{\theta}(\mathbf{x}).$ (6)
The parameterization is to be identified among a set of trajectories in phase space that describes the dynamics of the unknown system. Specifically, we denote the trajectory data by $\{(X_{i}^{j,0},X_{i}^{j,h})\}$, $1\leq i\leq N$, $1\leq j\leq M$, being $N$ the number of trajectories, $M$ the number of data pairs along each trajectory, and $X_{i}^{j,0}$ and $X_{i}^{j,h}$ the observed state positions for the $i$th trajectory at the time $t_{j}$ and $t_{j}+h$, respectively.
The networks $V_{\theta}$ and $\mathbf{g}_{\theta}$ are trained such that the error in the predicted dynamics is minimized and simultaneously to ensure the orthogonality condition $\mathbf{g}_{\theta}\cdot\nabla V_{\theta}=0$. Let $\mathcal{I}_{h}(X_{i}^{j,0};\mathbf{f}_{\theta})$ be the state at time $h$ for the system $\dot{\mathbf{x}}=\mathbf{f}_{\theta}(\mathbf{x})$ integrated by a suitable numerical scheme starting from $X_{i}^{j,0}$, then the loss function is taken as
$L(\theta)=$ $\frac{1}{NM}\sum_{i,j}\bigg{\lVert}\frac{1}{h}\Big{(}\mathcal{I}_{h}(X_{i}^{j,0};\mathbf{f}_{\theta})-X_{i}^{j,h}\Big{)}\bigg{\rVert}_{2}^{2}+\frac{\lambda}{S}\sum_{k=1}^{S}w\left(\frac{\nabla V_{\theta}(\tilde{X}_{k})\cdot\mathbf{g}_{\theta}(\tilde{X}_{k})}{|\nabla V_{\theta}(\tilde{X}_{k})|\cdot|\mathbf{g}_{\theta}(\tilde{X}_{k})|},\delta\right),$ (7)
in which the piecewise function $w(z,\delta)=z^{2}\mathbb{I}_{z\geq 0}+\delta z^{2}\mathbb{I}_{z<0}$ with $\delta=0.1$, $\lambda$ is a parameter controlling the strength of the loss term corresponding to the orthogonality condition, and $\{\tilde{X}_{k}\}$, $1\leq k\leq S$ is a representative subset of $\{X_{i}^{j,0}\}$ with a uniform distribution in the state space [21].
We parameterize the potential component $V(\mathbf{x})$ by a sum of a vanilla neural network $V^{\text{NN}}_{\theta}(\mathbf{x})$ with hyperbolic tangent ($\tanh$) as the activation function and a harmonic function, to make the learned system preserve good physical properties [21]. In addition, the training performances with the loss function (7) are ameliorated by adding: i) a normalization step to the input layer of $V_{\theta}$, $\mathbf{g}_{\theta}$ with respect to the data points; ii) a scalar multiplication to the output of the networks. In details, for the network $V_{\theta}$ we have
$V_{\theta}(\mathbf{x})$ $=\eta_{v}\cdot\tilde{V}_{\theta}\left(\frac{\mathbf{x}-\mu}{\sigma}\right),$ (8)
$\tilde{V}_{\theta}(\mathbf{y})$ $=V^{\text{NN}}_{\theta}(\mathbf{y})+|\mathbf{y}|^{2},$
with $\mu$ and $\sigma$ being the sample mean and deviation of the data points $\{X_{i}^{j,0}\}$, respectively. The parameter $\eta_{v}$ rescales the magnitude of $|\nabla V_{\theta}|$ to that of $|\mathbf{f}|$ estimated from the data. This is obtained by
$\eta_{v}$ $=\operatorname*{arg\,min}_{\eta>0}\sum_{i,j}\Big{(}\big{|}\eta\mathbf{v}_{i}^{j}\big{|}-\big{|}Y_{i}^{j}\big{|}\Big{)}^{2}$ (9)
$=\sum_{i,j}\big{|}\mathbf{v}_{i}^{j}\big{|}\cdot\big{|}Y_{i}^{j}\big{|}\bigg{/}\sum_{i,j}\big{|}\mathbf{v}_{i}^{j}\big{|}^{2},$
where the vector $\mathbf{v}_{i}^{j}=\nabla\tilde{V}_{\theta}\big{(}(X_{i}^{j,0}-\mu)/\sigma\big{)}$ and $Y^{j}_{i}=\big{(}X_{i}^{j,h}-X_{i}^{j,0}\big{)}/h$ is the estimated vector for $\mathbf{f}$ at the state $X_{i}^{j,0}$. Similarly, by estimating a multiplying scalar $\eta_{g}$ such that the magnitude of $|\mathbf{g}_{\theta}|$ approximates that of $|\mathbf{f}|$, the network $\mathbf{g}_{\theta}$ is
$\mathbf{g}_{\theta}(\mathbf{x})=\eta_{g}\cdot\mathbf{g}^{\text{NN}}_{\theta}\left(\frac{\mathbf{x}-\mu}{\sigma}\right).$ (10)
Note that the parameters $\mu$, $\sigma$, $\eta_{v}$ and $\eta_{g}$ are kept as constant during the entire training process.
2.3 Sparse symbolic regression of the quasipotential function
Based on the neural network approximations for the potential $V$ and circulatory vector $\mathbf{g}$, we further identify a sparse symbolic expression for the orthogonal decomposition of $\mathbf{f}$ that facilitates physical interpretation of the system. This is achieved by projecting the dynamics over a minimal set of functions from a candidate library $\bm{\Theta}(\mathbf{X})=[\theta_{1}(\mathbf{X}),\dots,\theta_{q}(\mathbf{X})]$, which is appropriately selected standing for possible functions. For $V$ and $\mathbf{g}$, the symbolic representation reads
$V^{\text{Symb}}(\mathbf{X})$ $=\bm{\Theta}(\mathbf{X})\ \bm{\Xi}_{v},$ (11)
$\mathbf{g}^{\text{Symb}}(\mathbf{X})$ $=\bm{\Theta}(\mathbf{X})\ \bm{\Xi}_{g},$
where $\bm{\Xi}_{v}$ and $\bm{\Xi}_{g}$ are the vectors containing the unknown coefficients to be determined. In addition, $\mathbf{X}$ represents a $NM$-by-$d$ matrix containing the data points $\{X_{i}^{j,0}\}$ from the sampled trajectories with multiple initial conditions. The domain of the symbolic regression does not necessarily coincides with that of the neural networks. Indeed, it is convenient not to overextend the identification far from the metastable states of the system. Outer boundaries of $V$ are inherently unreliable due to scarcity of data. We limit the selection over $\mathbf{X}$ by defining a subset of $\{X_{i}^{j,0}\}$ below a potential threshold value $\tau$, i.e. $V_{\theta}(\mathbf{x})<\tau$.
For a number of function libraries $\bm{\Theta}(\mathbf{X})$, e.g. polynomials or trigonometric functions, the gradient of the basis function, $\nabla\theta_{k}(\mathbf{X})$, can be expressed as a linear combination of the basis functions themselves. In this case, the gradient of the potential in Eq. (11) can still be represented via the same function library $\bm{\Theta}$, that is $\nabla V^{\text{Symb}}(\mathbf{X})=\bm{\Theta}(\mathbf{X})T(\bm{\Xi}_{v})$, where $T:\mathbb{R}^{q\times 1}\rightarrow\mathbb{R}^{q\times d}$ is a linear transformation function of the vector $\bm{\Xi}_{v}$. Therefore, the vector field is approximated as
$\mathbf{f}^{\text{Symb}}(\mathbf{X})$ $=-\nabla V^{\text{Symb}}(\mathbf{X})+\mathbf{g}^{\text{Symb}}(\mathbf{X})$ (12)
$=\bm{\Theta}(\mathbf{X})(-T(\bm{\Xi}_{v})+\bm{\Xi}_{g}).$
Now, existing sparse regression approaches aim to directly match observable data or byproducts such as their derivatives [29]. Unfortunately, directly inferring a symbolic regression of the quasipotential from data is not feasible. The essential information about the energy landscape remains concealed unless we possess prior knowledge on how to divide the dynamics into downhill and circulatory components. Therefore, we structure the symbolic identification to extract the quasipotential function from the decomposition provided by the neural networks. Specifically, a target matrix for the identification is set as
$\mathbf{G}(\mathbf{X})=[\mathbf{f}_{\theta}(\mathbf{X}),V_{\theta}(\mathbf{X}),\mathbf{g}_{\theta}(\mathbf{X})].$ (13)
We determine the unknown matrix $\bm{\Xi}\in\mathbb{R}^{q\times(2d+1)}$ which contains coefficient vectors for $\mathbf{f}^{\text{Symb}}$, $V^{\text{Symb}}$ and $\mathbf{g}^{\text{Symb}}$, via the constrained sparse relaxed regularized regression
$\operatorname*{arg\,min}_{\textbf{W},\bm{\Xi}}\,\frac{1}{2}\big{\lVert}\mathbf{G}(\mathbf{X})-\bm{\Theta}(\textbf{X})\,\bm{\Xi}\big{\rVert}^{2}+\lambda R(\textbf{W})+\frac{1}{2\nu}\big{\lVert}\bm{\Xi}-\textbf{W}\big{\rVert}^{2},$ (14)
where $R(\cdot)$ is a $l_{0}$-regularization term that promotes sparsity and minimizes over-fitting and W is an auxiliary variable which is introduced here to enable relaxation and partial minimization in order to improve the conditioning of the problem and tackle the non-convexity of the optimization [30]. In addition, $\lambda$ and $\nu$ are hyper-parameters that control the strength of regularization and relaxation, respectively. As indicated from Eq. (12),
$\bm{\Xi}=[-T(\bm{\Xi}_{v})+\bm{\Xi}_{g},\bm{\Xi}_{v},\bm{\Xi}_{g}]$ (15)
includes relations between the coefficients themselves. These are incorporated in the optimization problem (14) through constraints between the coefficients, denoted by
$\textbf{C}\ \bm{\Xi}=\textbf{0}.$ (16)
Remark. In general, the hypothesis space for $V$ or $\mathbf{g}$ may be chosen as the linear span of a subset of the library $\Theta(\mathbf{X})$, rather than utilizing the entire set. For example, one may perform regression in the space of polynomials in $\mathbf{x}$ up to fourth order for $V$ and polynomials up to third order for $\mathbf{g}$. The absence of some terms in a specific library subset can be readily achieved by adding constraining equations to Eq. (11). For instance, specific coefficients of $\bm{\Xi}_{v}$ and $\bm{\Xi}_{g}$ can be set to zero, which is incorporated into the optimization problem (14).
We solve the constrained optimization problem (14), (16) by performing an initialization step for the coefficient matrix $\bm{\Xi}$ (see A), obtaining $V^{\text{Symb}}$. The symbolic expression for the Freidlin-Wentzell quasipotential is its scalar multiple, $U^{\text{Symb}}(\mathbf{x})=2\cdot V^{\text{Symb}}(\mathbf{x})$.
3 Results
To provide a context for our study and to validate the efficacy of our machine learning approach, we investigate a system with a precisely known quasipotential (Sec. 3.1). Subsequently, we employ the combined data-driven technique to analyze the dynamics of a nanomechanical graphene resonator (Sec. 3.2).
In both the applications, we create a synthetic set of data points by sampling trajectories of the system and taking snapshots of the trajectories. In the first part of the method (Sec. 2.2), we train neural networks to reconstruct an orthogonal decomposition of the vector field. Details for sampling the data and training the neural networks are provided in B. In the second part of the method (Sec. 2.3), we identify sparse symbolic expressions for the downhill and circulatory components that define the dynamics of the system.
3.1 Example 1: An archetypal model with exact quasipotential
First, we consider a non-gradient dynamical system in the three-dimensional space,
$\left\{\begin{array}[]{l}\displaystyle\dot{x}=-2\left(x^{3}-x\right)-y-z\vspace{1mm}\\ \displaystyle\dot{y}=-y+2\left(x^{3}-x\right)\vspace{1mm}\\ \displaystyle\dot{z}=-z+2\left(x^{3}-x\right)\end{array}\right.$ (17)
The system has two stable equilibrium points at $A=(-1,0,0)$ and $B=(1,0,0)$ and the quasipotential with respect to the two states is given by $U(\mathbf{x})=x^{4}-2x^{2}+y^{2}+z^{2}+1$.
We generate $5000$ trajectories of the system by numerically simulating the dynamics (17) with initial states sampled from the computational domain $\Omega=[-2,2]\times[-1.5,1.5]\times[-1.5,1.5]$. Fig. 1 (a) shows a plot of five generated trajectories. A dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ is then constructed by taking snapshots from the generated trajectories. From the data, we first reconstruct an orthogonal decomposition of the vector field by training the neural networks $V_{\theta}(\mathbf{x})$ and $\mathbf{g}_{\theta}(\mathbf{x})$ with the loss function (7). Subsequently, we perform the sparse symbolic regression for the downhill and circulatory components in the decomposition. The identified components are showcased in Fig. 1 (b) and (c). Specifically, we take the basis functions for the regression as polynomials in $(x,y,z)$ up to fifth order for both $V^{\text{Symb}}$, $\mathbf{g}^{\text{Symb}}$. The target matrix is constructed with the neural network solutions over a subset of $\{X_{i}^{j,0}\}$ built with a potential threshold $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+2$. In the learning problem (14), we take the parameters $\lambda=0.1$ and $\nu=10^{-5}$.
Table 1: The exact and identified expressions for the quasipotential $U(\mathbf{x})$, downhill component $\nabla V(\mathbf{x})$ and circulatory component $\mathbf{g}(\mathbf{x})$ in the decomposition of the vector field for the system (17). Three digits after the decimal point are displayed in the identified expressions.
Exact expression       Identified expression
$U(\mathbf{x})$ $\quad x^{4}-2x^{2}+y^{2}+z^{2}+1$ $\hskip 14.22636pt1.000x^{4}-2.001x^{2}+0.999y^{2}$
$+0.999z^{2}+1.001$
$\nabla V(\mathbf{x})$ $\begin{bmatrix}2x^{3}-2x\\ y\\ z\end{bmatrix}$ $\hskip 14.22636pt\begin{bmatrix}2.001x^{3}-2.001x\\ 0.999y\\ 0.999z\end{bmatrix}$
$\mathbf{g}(\mathbf{x})$ $\begin{bmatrix}-y-z\\ 2x^{3}-2x\\ 2x^{3}-2x\end{bmatrix}$ $\hskip 14.22636pt\begin{bmatrix}-1.000y-1.000z\\ 2.003x^{3}-2.004x\\ 2.003x^{3}-2.003x\end{bmatrix}$
The identified quasipotential and downhill/circulatory components, $U^{\text{Symb}}(\mathbf{x})$, $\nabla V^{\text{Symb}}(\mathbf{x})$, $\mathbf{g}^{\text{Symb}}(\mathbf{x})$, as compared against the exact ones, are reported in Table 1. From the table, one can clearly observe that the identified expressions have the same like terms as in the exact ones with an error on the order of $10^{-3}$ for the corresponding coefficients. The results indicate that the proposed method is able to discover the explicit expression for the quasipotential of nonlinear systems from the observed data. Also, we show contour plots of the exact quasipotential $U(\mathbf{x})$ and identified potential $U^{\text{Symb}}(\mathbf{x})$ in Fig. 1 (d) and (e). The accuracy of the method is demonstrated by the fact that the learned potential $U^{\text{Symb}}$ is almost indistinguishable to $U$.
Figure 1: Combined data-driven method applied to the system (17) with $5000$ sampled trajectories. (a) Five generated trajectories. (b) The identified downhill component $-\nabla V^{\text{Symb}}(x,y,z)$. (c) The identified circulatory component $\mathbf{g}^{\text{Symb}}(x,y,z)$. (d) The exact quasipotential $U(x,y,z)$. (e) The identified quasipotential $U^{\text{Symb}}(x,y,z)$. All plots are projected on the $(x,y)$-plane. In panel (b)-(c), the line thickness shows the flow velocity.
Figure 2: (a) Phase plot of six observed trajectories and the trajectories generated by simulating the dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$ with the same initial states. (b)-(d) Plots of the invariant distribution computed by $p_{\epsilon}(x,y,z)=Z_{\epsilon}^{-1}\exp(-U^{\text{Symb}}(x,y,z)/\epsilon)$ of the randomly perturbed dynamics system (19) with various values of $\epsilon$ ($0.05$, $0.1$ and $0.2$). All plots are projected on the $(x,y)$-plane.
Long-term prediction. We proceed to assess the accuracy of the identified dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$ in predicting the evolution of the system. A set of 1000 trajectories ($\{X_{k}(t)\}$, $1\leq k\leq 1000$) not included in the previous ensemble are assessed. For each generated trajectory, we compute the error
$E_{k}=\frac{\big{\lVert}X^{\text{Symb}}_{k}(t)-X_{k}(t)\big{\rVert}_{2}}{\big{\lVert}X_{k}(t)\big{\rVert}_{2}}$ (18)
in order to quantify the difference between the original and learned dynamics. The statistics of the errors for the $1000$ generated trajectories are $3.04\times 10^{-4}\pm 6.55\times 10^{-3}$. The comparison of six generated trajectories with the observed ones in Fig. 2 (a) showcases the remarkable accuracy of the method.
Invariant distribution. With the identified quasipotential $U^{\text{Symb}}(\mathbf{x})$ as in Table 1, one is able to infer explicitly the invariant distribution of the system (17) in the presence of a white noise. The randomly perturbed dynamics is described by the stochastic differential equation:
$d\mathbf{x}_{t}=\mathbf{f}(\mathbf{x}_{t})dt+\sqrt{\epsilon}\ dW_{t},\quad t>0$ (19)
where $\epsilon$ is a parameter controlling the strength of the noise and $W_{t}$ is a Wiener process. When $\epsilon$ is small, the invariant distribution of the system (19) can be approximated via the quasipotential $U(\mathbf{x})$,
$p(\mathbf{x})=Z^{-1}\exp(-U(\mathbf{x})/\epsilon),$ (20)
where $Z$ is a normalization constant [12].
For the system (19), computing the invariant distribution $p_{\epsilon}(\mathbf{x})$ with the identified quasipotential $U^{\text{Symb}}(\mathbf{x})$ allows us to accurately estimate the normalization constant $Z$. This calculation is detailed in C. The estimated constants, by accounting for various values of $\epsilon$ ($0.05$, $0.1$ and $0.2$), are
$Z_{0.05}=0.0625;\quad Z_{0.1}=0.1794;\quad Z_{0.2}=0.5236.$ (21)
Therefore, the explicit expression of the invariant distribution is given by
$p_{\epsilon}(\mathbf{x})=Z_{\epsilon}^{-1}\exp\big{(}-U^{\text{Symb}}(\mathbf{x})/\epsilon\big{)}.$ (22)
Fig. 2 (b)-(d) illustrates the evolution of $p_{\epsilon}(\mathbf{x})$ for different values of $\epsilon$. As the noise intensity increases, the peaks of the invariant distribution broaden, with the probability density spreading out in the $(x,y)$ phase space.
3.2 Example 2: Nanomechanical graphene resonator
Here, the data-driven technique is applied to the dynamics of a graphene nanomechanical resonators. The equations
$\left\{\begin{array}[]{l}\displaystyle\dot{P}=\dfrac{{\omega_{0}}^{2}-{\omega_{F}}^{2}}{2\omega_{F}}Q-\zeta P+\dfrac{3}{8}\dfrac{\alpha}{\omega_{F}}Q\left(P^{2}+Q^{2}\right)\vspace{2mm}\\ \displaystyle\dot{Q}=-\dfrac{{\omega_{0}}^{2}-{\omega_{F}}^{2}}{2\omega_{F}}P-\zeta Q-\dfrac{3}{8}\dfrac{\alpha}{\omega_{F}}P\left(P^{2}+Q^{2}\right)-\dfrac{\beta}{2\omega_{F}}\text{.}\end{array}\right.$ (23)
describe the slow dynamics of a graphene membrane in terms of the $P$ and $Q$ variables that are the slowly varying in-phase and out-of-phase components of the motion, and obtained by applying a rotating wave approximation to the actual fast dynamics [6]. We take the following parameters in the dynamical equation: $\omega_{0}\!=\!1$, $\omega_{F}\!=\!1.0018$, $\zeta\!=\!0.00045$ and $\alpha\!=\!33$, and $\beta\!=\!1.4\!\times\!10^{-5}$. These describe a graphene drum in the classical duffing bistable region where both high and low amplitude oscillations occurs. The dynamics has two stable fix points at $A\approx(-0.007,-0.011)$ and $B\approx(0.004,-0.001)$, as shown in Fig. 3 (a).
Figure 3: (a) Five generated trajectories. (b) The observed dataset $\{X_{i}^{j,0}\}$, where the solid closed curve indicates the boundary of the region $\{\mathbf{x}\in\Omega:V_{\theta}(\mathbf{x})<\tau\}$ with a threshold value $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+6\times 10^{-9}$. (c) The dataset for the data matrix $\mathbf{X}$ sampled using the confined region (Right).
We create a synthetic dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ by generating $2000$ trajectories of the system (23) with initial conditions in a domain $\Omega=[-0.02,0.02]\times[-0.02,0.02]$. Five generated trajectories are shown in Fig. 3 (a). From the data, we learn an orthogonal decomposition of the vector field by training the neural networks $V_{\theta}(P,Q)$ and $\mathbf{g}_{\theta}(P,Q)$ with the loss function (7). Subsequently, we identify symbolic expressions for the components in the decomposition. A target matrix $\mathbf{G}(\mathbf{X})$ is constructed with the neural network solutions over a subset of $\{X_{i}^{j,0}\}$ built with a potential threshold $\tau=\min_{\Omega}V_{\theta}(\mathbf{x})+6\times 10^{-9}$. The full dataset used for training the neural networks, along with the reduced dataset utilized during the regression phase, are displayed in Fig. 3 (b) and (c). The basis functions in the identification are taken as polynomials in $(P,Q)$ up to fourth and third orders for $V^{\text{Symb}}$ and $\mathbf{g}^{\text{Symb}}$, respectively. We set the parameters as $\lambda=10^{-9}$ and $\nu=10^{-2}$ in the problem (14). An initial guess for the coefficient matrix $\bm{\Xi}_{v}$ and $\bm{\Xi}_{g}$ is obtained by performing regression for $V^{\text{Symb}}(\mathbf{x})$ and $\mathbf{g}^{\text{Symb}}(\mathbf{x})$ separately (see A).
Figure 4: Combined data-driven method applied to the system (23) with $2000$ sampled trajectories. (a) The identified downhill component $-\nabla V^{\text{Symb}}(P,Q)$. (b) The identified circulatory component $\mathbf{g}^{\text{Symb}}(P,Q)$. (c) The quasipotential $U(P,Q)$ estimated using a standard ordered upwind method [11]. (d) The potential $U_{\theta}=2V_{\theta}(P,Q)$ learned using the neural networks. (e) The potential $U^{\text{Symb}}(P,Q)$ identified using the sparse symbolic regression. In panel (a)-(b), the line thickness shows the flow velocity.
The identified symbolic expression (with rounded coefficients) for the quasipotential is
$U^{\text{Symb}}(P,Q)$ $=(8.12\times 10^{-9})-(1.56\times 10^{-6})P+(3.58\times 10^{-7})Q$ (24)
$\quad+(9.65\times 10^{-6})P^{2}+(1.71\times 10^{-5})PQ+(5.83\times 10^{-4})Q^{2}$
$\quad+0.023P^{3}+0.039P^{2}Q+0.047PQ^{2}+0.098Q^{3}$
$\quad+1.05P^{4}+1.88P^{3}Q+3.13P^{2}Q^{2}+2.47PQ^{3}+4.08Q^{4}.$
Also, the identified symbolic expression (with rounded coefficients) for the circulatory component, $\mathbf{g}^{\text{symb}}(P,Q)=[g_{1}^{\text{symb}}(P,Q),g_{2}^{\text{symb}}(P,Q)]$ is
$g_{1}^{\text{symb}}(P,Q)$ $=-(5.15\times 10^{-7})-(4.58\times 10^{-4})P-(1.88\times 10^{-3})Q$ (25)
$\quad+0.027P^{2}+0.048PQ-0.005Q^{2}$
$\quad+2.01P^{3}+14.61P^{2}Q+4.17PQ^{2}+11.99Q^{3},$
$g_{2}^{\text{symb}}(P,Q)$ $=-(6.61\times 10^{-6})+(1.77\times 10^{-3})P-(3.47\times 10^{-5})Q$
$\quad+0.012P^{2}+0.034PQ+0.116Q^{2}$
$\quad-11.60P^{3}+2.69P^{2}Q-9.68PQ^{2}+6.92Q^{3}.$
In Fig. 4 (a)-(b), we show plots of the identified downhill and circulatory components $\nabla V^{\text{Symb}}(P,Q)$, $\mathbf{g}^{\text{Symb}}(P,Q)$ in the decomposition of the vector field. Unlike the example in Sec. 3.1, we do not have knowledge of the true quasipotential. It is important to emphasize that this is the most common situation in real-world applications, and it further underlines the need of a data-driven approach. Thus, we find an alternative way to compare our estimation. Fig. 4 (c) reports a reference solution for the quasipotential $U$ using a standard ordered upwind method with a mesh of $1000\times 1000$ grid points [2]. The mesh-based method is compared with the identification provided by the neural network (Fig. 4 (d)), and the sparse symbolic regression (Fig. 4 (e)). From the figure, one can observe that the neural network (panel (d)) and symbolic (panel (e)) solutions agree well with the reference solution (panel (f)), which indicate the capability of the proposed method to infer the quasipotential from the observed data.
Figure 5: (a) Phase plot of six observed trajectories and the trajectories generated by simulating the dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$ with the same initial states. (b)-(d) Plots of the invariant distribution computed by $p_{\epsilon}(P,Q)=Z_{\epsilon}^{-1}\exp(-U^{\text{Symb}}(P,Q)/\epsilon)$ of the randomly perturbed dynamics system (19) with various values of $\epsilon$ ($10^{-9}$, $2\!\times\!10^{-9}$, and $3\!\times\!10^{-9}$).
Long-term prediction. Here we analyze the accuracy of the prediction for the identified dynamics $\dot{\mathbf{x}}=\mathbf{f}^{\text{Symb}}(\mathbf{x})$. Given $1000$ observed trajectories of the system (23) with initial states sampled from $\{\mathbf{x}\in\Omega:V_{\theta}(\mathbf{x})<\tau\}$, we generate trajectories by evolving the identified dynamics from the states. The statistics of the errors as in Eq. (18) for the generated trajectories are $0.0242\pm 0.0258$. Fig. 5 (a) shows a comparison of the generated trajectories with the observed ones, demonstrating the capacity to accurately capture the underlying dynamical evolution.
Invariant distribution. With the identified quasipotential $U^{\text{Symb}}(P,Q)$ of Eq. (24), we compute an explicit invariant distribution for the nanomechanical resonator modelled by  (23) when perturbed by a random noise. It is important to investigate how fluctuations affect the out-of-equilibrium dynamics of nonlinear resonators because atomically thin resonators are very sensitive to environmental noise. The invariant distribution of the randomly perturbed system is approximated by $p_{\epsilon}(P,Q)=Z_{\epsilon}^{-1}\exp(-U^{\text{Symb}}(P,Q)/\epsilon)$. Here the normalization constant is estimated by the integration over the domain $\Omega$. The evolution of the invariant distribution $p_{\epsilon}(P,Q)$ for different values of $\epsilon$, namely $10^{-9}$, $2\times 10^{-9}$ and $3\times 10^{-9}$, is showcased in Fig. 5 (b)-(d). While the predominance of the large oscillations of the nanomechanical resonator persists (motion around the large amplitude attractor - fixed point $A$), we observe a gradual smearing of the distribution in the phase space of the quadratures for larger $\epsilon$.
4 Conclusions
To conclude, we introduced a computational method for discovering explicit expressions of the quasipotential from observed data, which combines a neural network approximation and sparse regression. The efficacy of the method was illustrated using two model systems. The first example demonstrates the accuracy of the identified quasipotential, while the second example highlights the ability of the data-driven method to interpret nanomechanical vibrational systems of practical engineering interest.
A feature of the suggested approach is that it is directly applicable to both synthetic and experimental datasets since it only necessitates observing the nonlinear dynamics. The data-driven approach can be used to examine a wide variety of systems that undergo complex multistable behaviour, provided that sufficient data are available. Moreover, being based on neural networks and sparse regression makes the algorithm scalable to high-dimensional systems.
For the first time, explicit quasipotential expressions may be found without making any presumptions about the governing vector field. Because of this, we believe that the proposed approach is of interest to find a precise prediction of pertinent statistical values pertaining to the transition between metastable states. This new idea of determining a symbolic quasipotential function by coupling machine learning techniques offers a fully data-driven solution, and it demonstrates how recent developments in data science and machine learning open up intriguing opportunities in the realm of nonlinear dynamics.
Appendix A Initialization step for the problem of Eqs. (14) and (16)
An initial coefficient vector $\bm{\Xi}_{v}^{0}$ for the potential $V$ is obtained by
$\operatorname*{arg\,min}_{\textbf{W},\bm{\Xi}_{v}}\,\frac{1}{2}\big{\lVert}V_{\theta}(\mathbf{X})-\bm{\Theta}(\textbf{X})\,\bm{\Xi}_{v}\big{\rVert}^{2}+\lambda R(\textbf{W})+\frac{1}{2\nu}\big{\lVert}\bm{\Xi}-\textbf{W}\big{\rVert}^{2},$ (26)
where $V_{\theta}(\mathbf{X})$ is the target vector for $V$. Similarly, an initial coefficient matrix $\bm{\Xi}_{g}^{0}$ for the circulatory component $\mathbf{g}$ is produced by solving
$\operatorname*{arg\,min}_{\textbf{W},\bm{\Xi}_{g}}\,\frac{1}{2}\big{\lVert}\mathbf{g}_{\theta}(\mathbf{X})-\bm{\Theta}(\textbf{X})\,\bm{\Xi}_{g}\big{\rVert}^{2}+\lambda R(\textbf{W})+\frac{1}{2\nu}\big{\lVert}\bm{\Xi}-\textbf{W}\big{\rVert}^{2},$ (27)
where $\mathbf{g}_{\theta}(\mathbf{X})$ is the target matrix for $\mathbf{g}$. Therefore, with the relation (12) for the vector field $\mathbf{f}$, an initial value for $\bm{\Xi}$ for the problem (14) is given by
$\bm{\Xi}^{0}=[-T(\bm{\Xi}_{v}^{0})+\bm{\Xi}_{g}^{0},\bm{\Xi}_{v}^{0},\bm{\Xi}_{g}^{0}].$ (28)
Appendix B Data sampling and neural networks training
Table 2: The parameters for sampling data and training neural networks in the two numerical examples.
Parameters Archetypal model Nanomechanical graphene resonator
Sampling data $\Omega$ $[-2,2]\!\times\![-1.5,1.5]\!\times\![-1.5,1.5]$ $[-0.02,0.02]\!\times\![-0.02,0.02]$
$N$ $5000$ $2000$
$M$ $50$ $100$
$h$ $0.01$ $10$
$t_{j}$ $0.1j$ $100j$
Numerical integrator RK4 RK4
$V_{\theta}^{\text{NN}}$ Structure $3$-$50$-$50$-$50$-$1$ $2$-$100$-$100$-$100$-$1$
Activation $\tanh$ $\tanh$
$\mathbf{g}_{\theta}^{\text{NN}}$ Structure $3$-$50$-$50$-$50$-$3$ $2$-$100$-$100$-$100$-$2$
Activation $\tanh$ $\tanh$
Training Optimizer Adam Adam
Learning rate Exponentially decays Exponentially decays
Batch size $5000$ $5000$
Data points are collected from $N$ trajectories generated from numerical simulations of the dynamics. We employ a four-order Runge-Kutta integrator with time step $h/5$ and initial states sampled from a uniform distribution over the computational region $\Omega$. Then the dataset $\{X_{i}^{j,0},X_{i}^{j,h}\}$ is constructed by taking snapshots of the generated trajectories at times $t_{j}$ and $t_{j}+h$, $1\leq j\leq M$. In total, there are $5\times 10^{5}$ points in the dataset for the first example ($4\times 10^{5}$ points for the second example). We build a representative subset $\{\tilde{X}_{k}\}$ to impose the orthogonality condition as in Eq. (7) using the Algorithm 1 of Ref. [21] with a parameter $r$ of $0.1$ for the first example and $5\times 10^{-4}$ for the second example.
With the two datasets $\{X_{i}^{j,0},X_{i}^{j,h}\}$ and $\{\tilde{X}_{k}\}$, we train the neural networks $V_{\theta}(\mathbf{x})$ and $\mathbf{g}_{\theta}(\mathbf{x})$ using the loss function of Eq. (7). We utilize fully-connected neural networks for $V_{\theta}^{\text{NN}}$, $\mathbf{g}_{\theta}^{\text{NN}}$ with the activation function being the hyperbolic tangent ($\tanh$). In prior to training the networks, a normalization step and multiplying scalars are added to the input and output of the networks as in Eq. (8) and (10). In the loss function, we set the parameter $\lambda$ as $10$ and $10^{-11}$for the first and second examples, respectively, and take the numerical integrator $\mathcal{I}$ as the second order Runge-Kutta scheme. We train the neural networks for $1.5\times 10^{5}$ steps using the stochastic gradient descent with the Adam optimizer and a batch of $5000$ data points. In summary, the parameters used for sampling data and training neural networks are reported in Table 2.
Appendix C Estimation of the normalization constant in the invariant distribution for Example 1
We write the symbolic quasipotential $U^{\text{Symb}}(\mathbf{x})$ in Table 1 as
$U^{\text{Symb}}(\mathbf{x})=a_{1}x^{4}-a_{2}x^{2}+a_{3}+a_{4}y^{2}+a_{5}z^{2},$ (29)
where $a_{1}$, $a_{2}$, $a_{3}$, $a_{4}$ and $a_{5}$ are constant. The normalization constant in the invariant distribution $p_{\epsilon}(\mathbf{x})=Z_{\epsilon}^{-1}\exp\big{(}-U^{\text{Symb}}(\mathbf{x})/\epsilon\big{)}$ can be computed by
[…]
$=\sqrt{\frac{a_{2}}{8a_{1}}}\frac{\pi^{2}\epsilon}{\sqrt{a_{4}a_{5}}}\exp\left[\frac{a_{2}^{2}}{8a_{1}\epsilon}-\frac{a_{3}}{\epsilon}\right]\left[I_{1/4}\left(\frac{a_{2}^{2}}{8a_{1}\epsilon}\right)+I_{-1/4}\left(\frac{a_{2}^{2}}{8a_{1}\epsilon}\right)\right],$
where $I_{\alpha}(\cdot)$ denotes the modified Bessel function of the first kind [31].
Acknowledgments
The work of B. Lin was partially supported by A*STAR under its AME Programmatic programme: Explainable Physics-based AI for Engineering Modelling & Design (ePAI) [Award No. A20H5b0142]. PB acknowledges partial support from the European Union’s NextGenerationEU programme, in the framework of PRIN 2022, project DIMIN.
Data availability
Data supporting the findings in this manuscript are available upon request.
Competing interests
We declare we have no competing interests.